OpenStudy (anonymous):

Show how you solve the following two equations: x - 15 = 32 x - 7.2 = 23.1 I would really like some help please

OpenStudy (anonymous):

You'll get btter help in the proper section Take this to mathematics

OpenStudy (anonymous):

x - 15 = 32 +15 +15 x=32+15 x=47 x - 7.2 = 23.1 + 7.2 +7.2 x=23.1+7.2 x=30.3

OpenStudy (anonymous):

ffs I just said to go to maths.